OpenStudy (anonymous):

What mass of carbon is in a 245 g quantity of carbon dioxide (CO2)? A.3.75 g C B.66.8 g C C.81.7 g C D.1980 g C

OpenStudy (anonymous):

m(C) = m(CO2) * [Ar(C)/M(CO2)] = ???

OpenStudy (anonymous):

B 66.8 g C

OpenStudy (anonymous):

The molecular weight of CO2 is about 44 g/mol. Assume you have 1 mol of CO2. Therefore, there are 12 g C per 44 g CO2 (1 mol CO2). 12/44*245=66.8 g C